# Learning Frequency-Aware Dynamic Transformers for All-In-One Image Restoration

### v1

@@ IV-F Further analysis
Performance on spatially variant degradation. We analyze the performance of the proposed method under spatially variant degradation, aiming to highlight its enhanced capability in restoring spatial heterogeneity. Following the experimental setup of AirNet [15], we partition each clean image of the BSD68 [41] dataset into four regions. Subsequently, Gaussian noise with $\sigma\in\{0,15,25,50\}$ is injected into each region individually to create a new test set. We then assess the model trained solely on the standard denoising task using this new test set. As shown in Table VI, our method outperforms both AirNet [15] and PromptIR [18], achieving a PSNR improvements of 0.34 dB over AirNet and 0.11 dB over PromptIR.
@@
TABLE VI: Performance on spatially variant degradation. Under spatially variant degradation, our method showcases superior denoising performance compared to existing methods.
@@
TABLE VII: Generalization to unseen degradation level. Our method achieves superior generalization performance over the existing AirNet and PromptIR.
@@
Generalization to unseen degradation levels. To analyze the generalization capability of our model for unseen degradation levels, we evaluate it on the BSD68 [41] test set. Specifically, our model, trained solely on $\sigma\in\{15,25,50\}$, is tested with randomly sampled values from the ranges $\sigma\in[15,25]$ and $\sigma\in[25,50]$. The results shown in Table VII highlight the superior generalization performance of our model over AirNet and PromptIR.
The effect of the number of frequency bands. Finally, we examine the impact of the number of frequency bands, denoted as $L$, on both efficiency and performance. Specifically, we compare the performance between $L=2$ and $L=3$ under the standard ”noise-rain-haze” setting, as detailed in Table VIII. The results show that increasing the number of frequency bands from $L=2$ to $L=3$ improves restoration performance across all three tasks. This aligns with our expectations, as a higher value of $L$ enables the model to more finely distinguish between different degradation types at various frequencies, enhancing overall performance. However, the number of tokens in Intra- and Inter-Band attention scales with $L$, leading to attention maps and time complexity increasing proportionally. For example, the training time per epoch for Dformer is 70 seconds for $L=2$ and 90 seconds for $L=3$. To balance efficiency and performance, we have chosen to use $L=2$ as the default value for all experiments.
TABLE VIII: The effect of frequency decomposition on three degradation types. The restoration performance for all three tasks boosts when the number of frequency bands is increased from $L=2$ to $L=3$, while efficiency decreases
Method Denoise Derain Dehaze Training time of Dformer
@@ V Conclusion
This work presents an all-in-one image restoration model leveraging advanced vision transformers, inspired by the fact that various degradations uniquely impact image content across different frequency bands. The model consists of two primary components: the frequency-aware Degradation Prior Learning Transformer (Dformer) and the Degradation-Adaptive Restoration Transformer (Rformer). The Dformer captures degradation representations by using an input frequency decomposition module and frequency-aware Swin Transformer blocks. Guided by these learned representations, the Rformer utilizes a degradation-adaptive self-attention module to selectively focus on the most affected frequency components for restoration. Extensive experimental results demonstrate the superiority of our approach over existing methods in four key restoration tasks: denoising, deraining, dehazing, and deblurring. Furthermore, our method excels in handling spatially variant degradations and previously unseen degradation levels. These findings underscore the potential of our frequency-based perspective and advanced transformer design to significantly advance the field of image restoration.

### v2

@@ IV-G Generalization analysis
Performance on spatially variant degradation. We analyze the performance of the proposed method under spatially variant degradation, aiming to highlight its enhanced capability in restoring spatial heterogeneity. Following the experimental setup of AirNet [15], we partition each clean image of the BSD68 [68] dataset into four regions. Subsequently, Gaussian noise with $\sigma\in\{0,15,25,50\}$ is injected into each region individually to create a new test set. We then assess the model trained solely on the standard denoising task using this new test set. As shown in Table VIII, our method outperforms both AirNet [15] and PromptIR [17], achieving a PSNR improvements of 0.34 dB over AirNet and 0.11 dB over PromptIR.
@@
TABLE VIII: Performance on spatially variant degradation. Under spatially variant degradation, our method showcases superior denoising performance compared to existing methods.
@@
TABLE IX: Generalization to unseen degradation levels. Our method achieves superior generalization performance over the existing AirNet and PromptIR.
@@
TABLE X: Generalization to mixed degradations. Even with mixed degradations, our method maintains superior generalization performance.
Method Denoise + Derain Denoise + Deblur
Rain100L $(\sigma=25)$ GoPro $(\sigma=25)$
AirNet 29.17/0.858 22.61/0.667
PromptIR 29.41/0.857 22.50/0.665
Ours 30.02/0.862 25.03/0.728
Generalization to unseen degradation levels. To analyze the generalization capability of our model for unseen degradation levels, we evaluate it on the BSD68 [68] test set. Specifically, our model, trained solely on $\sigma\in\{15,25,50\}$, is tested with randomly sampled values from the ranges $\sigma\in[15,25]$ and $\sigma\in[25,50]$. The results shown in Table IX highlight the superior generalization performance of our model over AirNet and PromptIR.
Generalization to mixed degradations. To further analyze the generalization capability, we conducted experiments by mixing multiple degradations within the same image. Specifically, we injected Gaussian noise with a standard deviation of $\sigma$=25 to the degraded images from the rain removal dataset Rain100L [71] and the blur removal dataset GoPro [73], constructing a test set to assess the models trained under the 4 tasks All-in-one setting. The results, presented in Table X, demonstrate that our model exhibits superior generalization to mixed degradations.
The effect of the number of frequency bands. Finally, we examine the impact of the number of frequency bands, denoted as $L$, on both efficiency and performance. Specifically, we compare the performance between $L=2$ and $L=3$ under the standard “noise-rain-haze” setting, as detailed in Table XI. The results show that increasing the number of frequency bands from $L=2$ to $L=3$ improves restoration performance across all three tasks. This aligns with our expectations, as a higher value of $L$ enables the model to more finely distinguish between different degradation types at various frequencies, enhancing overall performance. However, the number of tokens in Intra- and Inter-Band attention scales with $L$, leading to attention maps and time complexity increasing proportionally. For example, the training time per epoch for Dformer is 70 seconds for $L=2$ and 90 seconds for $L=3$, with a corresponding increase in FLOPs from 213.35G to 231.10G. To balance efficiency and performance, we have chosen to use $L=2$ as the default value for all experiments.
TABLE XI: The effect of frequency decomposition on three degradation types. The restoration performance for all three tasks boosts when the number of frequency bands is increased from $L=2$ to $L=3$, while efficiency decreases
Method Denoise Derain Dehaze FLOPs Training time of Dformer
@@ V Conclusion
This work presents an all-in-one image restoration model leveraging advanced vision transformers, inspired by the fact that various degradations uniquely impact image content across different frequency bands. The model consists of two primary components: the frequency-aware Degradation Prior Learning Transformer (Dformer) and the Degradation-Adaptive Restoration Transformer (Rformer). The Dformer captures degradation representations by using an input frequency decomposition module and frequency-aware Swin Transformer blocks. Guided by these learned representations, the Rformer utilizes a degradation-adaptive self-attention module to selectively focus on the most affected frequency components for restoration. Extensive experimental results demonstrate the superiority of our approach over existing methods in five key restoration tasks: denoising, deraining, dehazing, deblurring, and low-light enhancement. Furthermore, our method excels in handling real-world degradations, spatially variant degradations, and previously unseen degradation levels. These findings underscore the potential of our frequency-based perspective and advanced transformer design to significantly advance the field of image restoration.